OpenStudy (anonymous):

–3x + 2y = 8 change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation.

OpenStudy (anonymous):

y-y1= m(x-x1) where (x1,y1) is the point.

OpenStudy (anonymous):

m is the slope

OpenStudy (anonymous):

Try to do it now.

OpenStudy (anonymous):

Idk..

OpenStudy (aroub):

slope intercept form: y=mx+b, where m is the slope and b is the y-intercept. => 2y=3x+8 y=3/2x+4 slope: 3/2 y-intercept: 4

OpenStudy (anonymous):

maybe you should learn now to do it.

OpenStudy (anonymous):

I am that's why I asked for help on here!